# NuiScene: Exploring Efficient Generation of Unbounded Outdoor Scenes

### v1

@@ 5 Conclusion
To establish a uniform scale across scenes, we first normalize all scenes to be within the $[-1,1]$ range. We then randomly select one scene as the anchor, assigning it a scale of 1. Each remaining scene is compared side by side with the anchor, and its scale is adjusted visually to align elements such as trees or buildings from multiple angles, to ensure a visually coherent relative size. Since these scenes are artist-created, their proportions may prioritize aesthetic appeal over real-world accuracy. As a result, there is no definitive correct scale. Instead, we approximate a visually consistent scaling and record the assigned scale for each scene, with the anchor scene remaining fixed at scale 1.
S1.3 Geometry Processing
@@
Point Cloud and Occupancy. All scene meshes are first converted into SDFs using the same method as Wang et al. [40], re-implemented in Taichi [16] for faster conversion. The scales obtained in S1.2 are used to adjust the voxel grid resolution for SDF conversion to enforce the unified scale across scenes. We then convert to occupancy by thresholding SDF values at zero. Finally, we apply multiple iterations of flood filling to fill in holes in the scene. Point clouds are sampled from the meshes after applying the Marching Cubes algorithm to the occupancy of each scene.
@@
S1.4 Chunk Sampling
To sample training chunks from the scenes, we first compute an alpha map from the top-down view of each scene. Next, we apply a convolution with all one kernel weights over the entire scene using a kernel size of $(100,100)$. The resulting convolution map is then thresholded at a value of $100\times 100$ to determine valid sampling locations. This ensures that all sampled chunks contain occupied regions, avoiding holes or boundary areas in the scenes.
Although we define the chunk size as $(50,50)$ along the x and z axes, we actually sample entire 2x2 chunks from the scene. This approach simplifies processing and better accommodates the diffusion model training. For the VAE model, the $(100,100)$ chunks are further divided into four $(50,50)$ chunks for training. An illustration of the sampling map is provided in Figure 9.
Additionally, we compute a depth variation map. For each pixel, we calculate the mean depth within the kernel window and subtract the pixel’s depth value, taking the absolute value to obtain depth mean deviations. To avoid sampling overly flat regions, we filter out sample locations where the depth variation is below 2.5.
@@
S1.5 NuiScene43 Statistics
We present statistics of NuiScene43 in Tab. 6, including the number of scenes in each category and the total number of quad chunks that can be sampled from each category. The chunk count is derived by summing the sampling map discussed earlier, following depth variation thresholding. Note that these chunks may overlap, as we consider all valid x and z coordinates for sampling.
@@
1:Number of rows $I$, number of columns $J$, trained diffusion model $\epsilon_{\theta}$
@@
3:for $i=0$ to $I-2$ do $\triangleright$ Iterate row-wise
4:       for $j=0$ to $J-2$ do $\triangleright$ Iterate column-wise
@@
18:             $\bm{X}_{t}\sim\mathcal{N}(0,\mathbf{I})$triangleright Sample Gaussian noise
19:             $\bm{C}\leftarrow M\oplus Z_{\text{cond}}\oplus\text{PE}$triangleright Conditioning input
20:             $\{\bm{z}_{0},\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}\}\leftarrow\epsilon_{\theta}(\bm{X}_{t}\oplus\bm{C},t)$triangleright Diffuse 2×2 quad chunk
21:             $\bm{Z}_{i,j},\bm{Z}_{i,j+1},\bm{Z}_{i+1,j},\bm{Z}_{i+1,j+1}\leftarrow\bm{z}_{0},\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}$triangleright Write to scene latent grid
@@
Appendix S2 Implementation Details
S2.1 Raster Scan Order Generation
We show the algorithm of our raster scan order generation during inference utilizing our diffusion model trained with 4 different configurations in Algorithm 1. The algorithm outlines how the 4 different masking configurations are used based on different positions to condition on existing chunks and enable continuous and unbounded generation. After we obtain a large grid of chunks embeddings with Algorithm 1, the decoder of the VAE is utilized to decode occupancies for the scene chunks. We then use marching cubes to convert the large scene grid of occupancies into the final mesh.
@@
S2.2 Multi-scene training
For training the multi-scene model in the paper we add an additional 3 scenes along with the original single scene for training. The additional 3 scenes are shown in Figure 10.
S2.3 Details on Scene Texturing
@@
SceneTex [6] offers three camera modes [^1] https://github.com/daveredrum/SceneTex for rendering images and texture optimization, i.e., Spherical Cameras, the Blender Cameras and BlenderProc Cameras. For small indoor scenes like those from 3DFront, spherical cameras are sufficient to capture the fine-grained details of the scene and produce high-quality textures. However, for large outdoor scenes, a predefined spherical camera trajectory often fails to cover the entire scene comprehensively and misses important scene details. We therefore choose the Blender Cameras mode to manually keyframe the camera trajectory for each scene. This enables a finer control over capturing the scene details, leading to an improved texture.
@@
More specifically, we first normalize each scene into a spatial range of $[-1,1]^{3}$ centered at the world origin. Then, we define a snake-scan-like camera trajectory to capture the entire scene at a specified frame rate, as shown in Figure 11. Each frame will be rendered by SceneTex and all rendered images are sampled during training to optimize the scene-level texture. Note that to enable faster optimization, we use Blender’s Decimate Geometry and Merge By Distance with a distance of 0.001m to compress the scene until its size is smaller than 20MB. This might degrade the geometry for large scenes, but it’s only for the texturing purpose. We encourage readers to focus on the untextured geometry details of the generated scenes.
Figure 12 shows more examples on raw geometries and the textured counterparts.
Appendix S3 Additional Results
S3.1 Single Scene Results
We show additional results from our diffusion model trained on a single scene with vector set backbone in Figure 13. And additional results for the RePaint baseline in Figure 14.
S3.1.1 Multi Scene Results
We show additional results from our diffusion model trained on 4 scenes with vector set backbone in Figure 15 and Figure 16.
@@
Figure 15: Additional results for our vector set diffusion model trained on multiple scenes. The scenes shown here are of size $21\times 21$.
@@
Figure 16: Additional results for our vector set diffusion model trained on a 4 scenes. The scenes shown here are of size $16\times 46$. The scene in the middle is the untextured mesh of Figure 2.

### v2

@@ A.2 Scale Labeling
To establish a uniform scale across scenes, we first normalize all scenes to [1,1][-1,1]. We then randomly select one scene as the anchor, assigning it a scale of 1. Each remaining scene is compared with the anchor, and its scale is adjusted visually to align elements such as trees or buildings from multiple angles, to ensure a visually coherent relative size. Since these scenes are artist-created, their proportions may prioritize aesthetic appeal over real-world accuracy. As a result, there is no definitive correct scale. Instead, we approximate a visually consistent scaling and record the assigned scale for each scene, with the anchor scene remaining fixed at scale 1.
A.3 Geometry Processing
@@
Point Cloud and Occupancy. All scene meshes are first converted into SDFs using the same method as Wang et al. [43], re-implemented in Taichi [17] for faster conversion. The scales obtained in A.2 are used to adjust the voxel grid resolution for SDF conversion to enforce the unified scale across scenes. We then convert to occupancy by thresholding SDF values. Finally, multiple iterations of flood filling are applied to fill in holes in the scene. Point clouds are sampled from the meshes after applying the Marching Cubes algorithm to the occupancy of each scene.
@@
A.4 Chunk Sampling
To sample training chunks from scenes, we first compute an alpha map from the top-down view. Next, we apply a convolution with all one kernel weights over the entire scene using a kernel size of $(100,100)$, the size of a quad chunk along the $x$ and $z$ axis. The resulting convolution map is then thresholded at a value of $100\times 100$ to determine valid sampling locations. This ensures that all sampled chunks contain occupied regions, avoiding holes or boundary areas in the scenes.
Next, we compute a depth variation map. For each pixel, we calculate the mean depth within the kernel window and subtract the pixel’s depth value, taking the absolute value to obtain depth mean deviations. To avoid sampling overly flat regions, we filter out sample locations where the depth variation is below 2.5. An illustration of the sample maps is provided in Figure˜10.
@@ A.4 Chunk Sampling
A.5 NuiScene43 Statistics
We present statistics of NuiScene43 in Tab.˜6, including the number of scenes in each category and the total number of quad chunks that can be sampled. The chunk count is derived by summing the sampling map discussed earlier, following depth variation thresholding. Note that these chunks may overlap, as we consider all valid $x$ and $z$ coordinates.
@@ A.5 NuiScene43 Statistics
1:Number of rows II, number of columns $J$, trained diffusion model $\epsilon_{\theta}$
@@
3:for $i=0$ to I2I-2 do $\triangleright$ Iterate row-wise
4:   for $j=0$ to J2J-2 do $\triangleright$ Iterate column-wise
@@
18:     $\bm{X}_{T}\sim\mathcal{N}(0,\mathbf{I})$ $\triangleright$ Sample Gaussian noise
19:     $\bm{C}\leftarrow M\oplus Z_{\text{cond}}\oplus\text{PE}$ $\triangleright$ Conditioning input
20:     $\{\bm{z}_{0},\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}\}\leftarrow\text{Denoise}(\bm{X}_{T}\oplus\bm{C},\epsilon_{\theta})$ $\triangleright$ Denoise 2×2 quad chunk
21:     $\bm{Z}_{i,j},\bm{Z}_{i,j+1},\bm{Z}_{i+1,j},\bm{Z}_{i+1,j+1}\leftarrow\bm{z}_{0},\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}$ $\triangleright$ Write to scene latent grid
@@
Appendix B Implementation Details
B.1 4-scene Training
We add an additional 3 scenes along with the original single scene for training the 4-scene model in the main paper. The 3 additional scenes are shown in Figure˜11.
B.2 Raster Scan Order Generation
We show the algorithm of our raster scan order generation during inference utilizing our diffusion model trained with 4 different configurations in Algorithm˜1. The algorithm outlines how the 4 different masking configurations are used to condition on existing chunks for continuous and unbounded generation. In line 19 of the algorithm we denoise for 50 steps using the DDPM scheduler. After we obtain a large grid of embeddings, the decoder is used to decode occupanices followed by marching cubes to get the final mesh.
It can be observed that generation of the next row can begin once a single quad chunk has been generated in the current row. This allows for parallelization along the scene’s anti-diagonal, greatly accelerating sampling for large scenes. However, memory usage becomes inconsistent and varies with scene size. For consistency and simplicity, we use the sequential raster scan order (Algorithm˜1) in the paper. Please refer to our implementation for more details.
B.3 Details on Scene Texturing
@@
SceneTex [6] offers three camera modes [^2] https://github.com/daveredrum/SceneTex for rendering images and texture optimization, i.e., Spherical Cameras, the Blender Cameras and BlenderProc Cameras. For small indoor scenes like those from 3DFront [12], spherical cameras are sufficient to capture the fine-grained details of the scene and produce high-quality textures. However, for large outdoor scenes, a predefined spherical camera trajectory often fails to cover the entire scene comprehensively and misses important scene details. We therefore choose the Blender Cameras mode to manually keyframe the camera trajectory for each scene. This enables a finer control over capturing the scene details, leading to an improved texture.
@@
More specifically, we first normalize each scene into a spatial range of $1,1]^{3}$ centered at the world origin. Then, we define a snake-scan-like camera trajectory to capture the entire scene at a specified frame rate, as shown in Figure˜12. Each frame will be rendered by SceneTex and all rendered images are sampled during training to optimize the scene-level texture. Note that to enable faster optimization, we use Blender’s Decimate Geometry and Merge By Distance with a distance of 0.001m to compress the scene until its size is smaller than 20MB. This might degrade the geometry for large scenes, but it’s only for the texturing purpose. We encourage readers to focus on the untextured geometry details of the generated scenes. Figure˜13 shows more examples on raw geometries and the textured counterparts.
Appendix C Additional Results
C.1 Single Scene Results
Table 7: Additional quantitative comparison of reconstruction across different VAE backbones for the single scene experiment. Here $\hat{h}$ indicates that the predicted height was used for the occupancy prediction and $\tilde{h}$ the ground truth height.
[…]
triplane $3\times 64^{2}$/66 0.940 0.064 0.831 0.064 0.831
$3\times 128^{2}$/66 0.982 0.185 $\mathbf{0.879}$ 0.058 0.858
vecset - $\mathbf{0.989}$ $\mathbf{0.055}$ 0.864 $\mathbf{0.055}$ $\mathbf{0.863}$
Triplane $3\times 128^{2}$. Tab.˜7 presents results for the Triplane baseline with an output resolution of $3\times 128^{2}$. We see that the IOU outperforms the lower-resolution triplane baseline and beats the vecset model on F-Score($\hat{h}$). However, we notice a large increase in CD($\hat{h}$). Upon inspecting chunks with larger CD values, when predicted heights are larger than the ground truth, the model produced floating artifacts. This may be attributed to the higher resolution, which may be more sensitive to spatial aliasing or quantization artifacts during feature querying. Adjusting the scale factor to $S=8.5$ may be more appropriate for this resolution, but due to the high training cost, we did not retrain the model. These results highlight the sensitivity of spatially structured latents to scene bounds and hyperparameters like $S$.
Table 8: Comparison of VAE training resources for vector set vs triplane backbones with larger batch sizes. Training for all experiments was run on 4 L40S GPUs, total batch size and memory across 4 gpus are reported. The # Latents is the size of the latent for the VAE backbone and Output Res indicates the triplane size after deconvolution.
Method BS # Latents Output Res Time (hr) Mem. (GB)
triplane 192 $\times 4^{2} 3×3$$\times 32^{2}$ 18 166.0
120 $3\times 4^{2}$ $3\times 64^{2}$ 29.4 164.4
vecset 144 16 - 21.6 170.5
Table 9: Quantitative comparison of reconstruction across different VAE backbones using larger batch sizes in Tab.˜8. Here $\hat{h}$ indicates that the predicted height was used for the occupancy prediction and $\tilde{h}$ the ground truth height.
[…]
triplane $3\times 32^{2}$/66 0.727 0.171 0.508 0.170 0.503
$3\times 64^{2}$/66 0.933 0.099 0.852 0.064 0.830
vecset - $\mathbf{0.962}$ $\mathbf{0.072}$ $\mathbf{0.890}$ $\mathbf{0.057}$ $\mathbf{0.858}$
Figure 14: We compare the reconstruction of the vector set model with larger batch size (144) when using the predicted height $\hat{h}$ of the model and using the ground truth height $\tilde{h}$.
Figure 15: Here we show the results of using different occupancy grid resolutions for prediction during inference for marching cubes. The numbers on the bottom indicate the chunk size along the $x$ and $z$ axis. $50$ is the original chunk size used for training.
@@ C.1 Single Scene Results
Larger Batch Size. We initially experimented with larger batch sizes across all configurations, using 4 GPUs, as shown in Tab.˜8 to accelerate training. For the vecset model, this reduced training time to 21.6 hours from 36.1 hours on 2 GPUs (main paper configuration). However, it led to overfitting, as seen in Tab.˜9 which also affected height prediction accuracy as demonstrated in Figure˜14 where the model under-predicted the height, leaving parts of the chunk unreconstructed. In contrast, triplane backbones were less affected by overfitting. We hypothesize this is due to their spatial locality where query points retrieve features from fixed triplane positions, whereas the vecset backbone allows queries to aggregate from all feature vectors, offering greater capacity but also more prune to overfitting. Nevertheless, this shows the limit of the one scene training.
Evaluation Metrics. Following 3DShape2Vecset [52], we report IOU, Chamfer Distance (CD), and F-Score for VAE reconstruction. However, we find that CD and F-Score can be less reliable. When extracting surfaces via marching cubes, a level set threshold must be selected, which can introduce gaps relative to the ground truth due to the discrete grid resolution. This issue can disproportionately affect chunks of varying height, taller chunks may have fewer points per area given the same number of points are sampled for each chunk, leading to potentially higher CD and lower F-Score. As a result, these metrics are not fully comparable across settings like single-scene versus 4-scene training, due to differences in chunk distributions. In our case, larger differences in reported values are more indicative, while smaller variations may reflect noise due to the level set gap. Overall, IOU is more reliable and comparable across different scene configurations.
Level of Detail. In Figure˜15, we visualize occupancy predictions at different resolutions during inference with marching cubes to generate different level of detail (LoD) for the triplane and vector set models. $50$ is the original chunk size used for training along the $x$ and $z$ axis, with the $y$ axis scaled according to the height prediction. We find that halving the chunk size offers comparable details while reducing memory usage for scenes. However, decreasing further to $13$ leads to noticeable loss of details in the trees and buildings.
Additional Visualizations We show additional results for our vecset diffusion model in Figure˜16 and results for the RePaint baseline in Figure˜17.
@@
Figure 18: Additional results for our vector set diffusion model trained on 4 scenes. The scenes shown here are of size $21\times 21$.
@@
Figure 19: Additional results for our vector set diffusion model trained on 13 scenes. The scenes shown here are of size $21\times 21$.
C.2 4-scene
Table 10: Quantitative comparison of reconstruction across different VAE backbones on the 4 scene training. Here $\hat{h}$ indicates that the predicted height was used for the occupancy prediction and $\tilde{h}$ the ground truth height.
[…]
triplane $3\times 64^{2}$/66 0.933 0.061 0.833 0.061 0.832
vecset - $\mathbf{0.989}$ $\mathbf{0.053}$ $\mathbf{0.869}$ $\mathbf{0.053}$ $\mathbf{0.868}$
Table 11: Comparison of triplane and vecset diffusion models for generated quad-chunks on the 4 scene training. KPD scores are multiplied $10^{3}$.
[…]
triplane 0.868 2.142
vecset 0.581 1.104
For the 4-scene experiments in this section we follow the same GPU configurations in the main paper, and increase the number of sampled chunks across 4 scenes to 300K chunks. The training and validation follow the same 95%-5% split. For diffusion evaluation we sample 30K quad chunks from the scenes and well as the diffusion models for evaluation.
Quantitative Results. We show the VAE reconstruction and diffusion quality in Tab.˜10 and Tab.˜11, respectively. We can see that for IOU the vector set model maintains its performance, while the triplane slightly drops. With the vector outperforming the triplane model across the board.
Additional Visualizations. We show additional results from our vecset diffusion model trained on 4 scenes in Figure˜18.
C.3 13-scene
To further demonstrate our model’s potential for scaling up, we sample 280K chunks from 13 scenes in NuiScene43 (sample configurations available here) and train using the same settings as before.
Improving Reconstruction. Following the previous vector set settings, the VAE model trained on 13 scenes see a drop in IOU to $0.968$, due to a larger variation of scenes compared to the single or 4-scene scenarios. One naive way we can increase the VAE’s performance is to increase the number of feature vectors directly during training from $V=16$. However, this would increase the memory usage and slow diffusion training as shown in the triplane model. Instead, inspired by the deconvolution layers for triplanes. We introduce a pixel shuffle [40]-like layer for upsampling the number of vector sets followed by additional self attention layers. This increases the capacity of the model without increasing the latent size. Specifically, we add an additional projection layer that increases the number of channels and reshape from the channels dimension to the vector tokens. We upsample to $512$ vector sets and add 3 self-attention layers before the cross attention layer for querying coordinates. This improves the IOU to $0.983$.
In Figure˜20, we show a qualitative comparison between models. The original vector set model already reconstructs scenes quite well, despite the increase in the number of training scenes. The added improvements further enhance the reconstruction of finer details.
Figure 20: We compare chunk reconstructions between the original vector set model and the upsampled version (VecSet + PS). The upsampling model recovers finer details that are blurry in the original (see orange boxes).
Additional Visualizations. We show square scenes generated by the diffusion model trained with the improved model with pixel shuffle in Figure˜19. And larger scenes with aerial and street view zoom-ins in Figure˜9. It can be seen that this model can generate a larger variety of different scenes while maintaining good fidelity.